OpenStudy (anonymous):

simplify 7divided by 2over 7 explain

OpenStudy (anonymous):

\[\frac{7 \div 2}{7}\]\[=\frac{\frac{7}{2}}{7}\] With me so far?

OpenStudy (anonymous):

Hrm.. I take that as a no.